OpenStudy (anonymous):

Please help? Look for equation in Comments. Question: Find the equation of the circle whose center and radius are given.

OpenStudy (anonymous):

\[center ( 7, -3), radius =\sqrt{7}\]

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2 (x-7)^2+(y+3)^2=49 x^2+y^2-14x+6y+7=0 is the equation